Hello again im having trouble with this one too. help please.
/ (fraction sign)
1/2x-1/8y=-1/4
-4x+y=2

Found 2 solutions by RAY100, stanbon:
Answer by RAY100(1637)

(1/2)x - (1/8)y = -(1/4),,,,,mult by 8
.
4x -y=-2,,,,,,,,mult by -1
.
-4x +y = 2
.
But this matches the second equation, therefore the equations are coincident(lay of top of each other).
.
Therefore there are infinitely many solutions
.

Answer by stanbon(75887)

Get rid of the fractions:
-------------------
(1/2)x-(1/8)y=-1/4
-4x+y=2
---------------------------
4x - y = -2
-4x+ y = 2
----------------
Add those and you will get:
0 = 0
===================================
That means the two equations are really one.
The solution is any point that fits
into (x, 4x+2), for any value of x.
============================================
Cheers,
Stan H.